# The LHAASO Collaboration

### v2

@@ I Introduction
Binary systems are prominent sources in the high-energy sky above a few keV, with X-ray binaries being the dominant population [1]. However, only a small number of these systems have been observed at more energetic gamma-ray bands [2, 3]. Among them, a distinct class known as gamma-ray binaries has been identified, characterized by persistent gamma-ray emission and spectral energy distributions ($\nu F_{\nu}$) that peak in the gamma-ray range [2, 4]. All known gamma-ray binaries consist of a compact object in orbit around a massive, young O- or Be-type star, with clear orbital modulation observed across multiple wavelengths [2]. These modulations indicate that the gamma-ray emission is likely powered by intra-binary interactions — such as collisions between a relativistic outflow from the compact object and the dense stellar wind or radiation field of the companion star. TeV emission was detected from these compact systems, but it is unclear about the maximal energy reached by binary systems and whether they are capable of accelerating protons to PeV energies. Probing these systems at ultra-high-energy (UHE; >>100 TeV) gamma-rays may open a unique window to study extreme particle acceleration, relativistic outflows, and the accretion–ejection processes operating in compact binary environments.
LS I +61 303 is a prototypical gamma-ray binary system composed of a massive Be-type star and a compact object of uncertain nature [5, 6]. Recent reports of transient radio pulsations from the direction of LS I +61 303 strongly suggest that the compact object is a rotating neutron star [7]. Located at a distance of $\sim$2 kiloparsecs [8], LS I +61 303 exhibits a well-established orbital period of 26.496(8) days [9], confirmed across optical [10], X-ray [11, 12], and gamma-ray [13] bands. In 2006, the MAGIC telescope detected variable TeV gamma-ray emission modulated by the orbital period, with the short variability time scale suggesting a compact emission origin [14]. This orbital modulation was subsequently confirmed by VERITAS [15]. Both instruments detected enhanced TeV emission near apastron.
The Large High Altitude Air Shower Observatory (LHAASO) is a large hybrid observatory designed to explore the gamma-ray sky from $\sim 100$ GeV to several PeV [16]. It consists of three major detector arrays: the Kilometer Square Array (KM2A), the Water Cherenkov Detector Array (WCDA) and the Wide Field-of-View Cherenkov Telescope Array (WFCTA) [17, 18, 19]. In this paper, we present LHAASO observations of LS I +61 303, spanning roughly 1 TeV to several hundred TeV. These results provide unprecedented access to the ultra-high-energy regime and offer new constraints on the nature of particle acceleration in gamma-ray binaries.
@@ II Methods
Data from both WCDA and KM2A are used in this analysis. The WCDA data are collected with the full array from March 2021 to July 2024, and the number of triggered detectors($N_{\rm{hit}}$) serves as a proxy for energy binning. The pincness parameter is used to separate gamma-rays from the cosmic ray background. The residual cosmic ray background after $\gamma$/hadron separation is estimated using the “direct integration method”.
The KM2A data used here are collected from December 2019 to July 2024, yielding an effective exposure of 1570.4 days after selection. The event-reconstruction algorithm and selection criteria are detailed in the performance paper [19]. The data are binned into five bins per energy decade based on the reconstructed energy, and the cosmic-ray background is estimated with the “direct integration method”.
A 3D likelihood fitting process is used to determine the spectrum and morphology simultaneously. The likelihood ratio test compares the goodness of fit between hypotheses. The test statistic is defined as ${\rm TS}=2\ln(\lambda)$, with $\lambda=\mathcal{L}_{s+b}/\mathcal{L}_{b}$, where $\mathcal{L}_{s+b}$ is the maximum likelihood for the source-plus-background model and $\mathcal{L}_{b}$ is that for the background-only model. Following Wilks’s theorem[20], TS is converted to significance: under the null hypothesis TS follows a $\chi^{2}$ distribution whose degrees of freedom equal the difference of free parameters between the background-only and source-plus-background models.
@@
To maintain consistent detector performance, only data taken with the full array from July 2021 to July 2024 are used for the orbital-variability analysis. The orbital phase of each event is calculated as $\phi=((T-T_{0})~\bmod~P)/P$, where ${\rm T}_{0}={\rm MJD}~43366.275$ is the zero phase epoch [21], ${\rm P}=26.496$ d is the orbital period [9], and mod denotes the modulo operator. We adopt the widely accepted periastron phase $\phi_{\rm peri}=0.23$ [22, 23], although some alternative orbital solutions yield a slightly later phase ($\phi_{\rm peri}=0.275$; [24]). The data are divided into ten phase bins per orbit. A likelihood-ratio test is employed: the null hypothesis assumes constant flux in all bins, whereas the alternative hypothesis allows the flux to vary among bins. Under Wilks’s theorem, the resulting TS follows a $\chi^{2}$ distribution with 9 degrees of freedoms.
Figure 1: Test-statistic maps after subtraction of the contributions from nearby sources and Galactic diffuse emission around LS I +61 303, observed with WCDA (A) and KM2A (B), for the energy ranges 1.5–30.5 TeV and >>25 TeV, respectively. Red crosses indicate the position of LS I +61 303 at optical band [25], while the green ‘x’ represent the position of LS I +61 303 as detected by LHAASO. The black dashed line indicates the positional uncertainty of LS I +61 303 at the 95% level, which is consistent with the position measurement at X-ray band. PSF at each energy band is marked by double-headed arrow.
[…]
Figure 3: (A,B) Flux measured by WCDA (1.4–30.5 TeV) and KM2A (25–100 TeV) versus orbital phase. (C) Energy of photons >100>100 TeV measured by KM2A versus orbital phase. Two orbital cycles are shown for clarity.
@@ III Results
The gamma-ray spectrum shown in Fig. 2 can be described ($\chi^{2}/ndf=6.7/8$) by a single power law model $d{\rm N}/d{\rm E}=N_{0}\times(E/10{\rm TeV})^{-\Gamma}$, where $N_{0}$ = $(2.18\pm 0.20)\times 10^{-15}~\rm{TeV}^{-1}~{cm}^{-2}~{s}^{-1}$ is the flux normalization and $\Gamma$$3.00\pm 0.05$ is the photon index. The integral flux above 1 TeV, $F_{>1{\rm TeV}}=1.10\times 10^{-12}~\rm{cm}^{-2}~{s}^{-1}$, corresponds to $3.8\%$ of the Crab unit. The quoted errors are statistical only. Systematic uncertainties are assessed separately for WCDA and KM2A using the Crab Nebula as a standard candle, yielding $^{+8}_{-24}\%$ and $\pm 7\%$ for flux measurements, respectively. The difference between the atmospheric model used in the simulation and the actual atmosphere dominates the systematic error for KM2A; this introduces an uncertainty of 0.02 in the spectral index. Fig. 2 also shows the spectra measured by VERITAS and MAGIC for comparison. Although the flux of this source shows obvious orbital modulation, and the orbitally-averaged flux has shown slight changes on long time-scales (i.e. super-orbital modulation; [28, 23]), the TeV gamma-ray flux here is roughly consistent with those measured by MAGIC [14, 29] and VERITAS [26].
@@ IV Discussion and conclusion
In this study, we report the first definite detection of gamma-ray emission up to UHE range from the gamma-ray binary LS I +61 303 using LHAASO observations. The extension of the spectral energy distribution to energies nearly an order of magnitude higher than previous measurements (e.g., by MAGIC and VERITAS) not only challenges conventional acceleration models but also provides a unique opportunity to probe the balance between particle acceleration and energy losses in such systems.
@@
In the leptonic scenario, high-energy electrons upscatter stellar photons via inverse Compton (IC) scattering. However, in the UHE regime, where the IC scattering enters the Klein–Nishina regime (For a 100 TeV electron, $\gamma\varepsilon\gg m_{e}c^{2}$ with $\varepsilon=1.9$ eV, corresponding to a Be-star temperature of 22.5 kK [30]), the IC process will be largely depressed. The IC process is also constrained by the maximum energy of electrons can be accelerated in binary systems, which is fundamentally determined by the the particle acceleration efficiency, cooling time and the limited size. Early studies [31, 32] have shown that the maximum energy is sensitive to parameters such as the magnetic field strength, the efficiency of the acceleration mechanism, and the ambient photon density. Assuming an acceleration timescale $\tau_{\rm acc}=E/(\eta eBc)\approx 1.1(E/1{\rm TeV})(\eta/0.1)^{-1}(B/1{\rm G})^{-1}$ s with efficiency factor $\eta\approx 0.1$ and a magnetic field $B\approx 0.2$ G [33], and a synchrotron cooling timescale $t_{\rm syn}=6\pi m_{e}^{2}c^{3}/(\sigma_{T}B^{2}E)\approx 400(E/1{\rm TeV})^{-1}(B/1{\rm G})^{-2}$ s, the maximum energy $E_{\rm max}$ can be obtained by equating $\tau_{\rm acc}$ to the synchrotron cooling timescale $t_{\rm syn}$, then $E_{\rm max}=\sqrt{6\pi\eta em_{\rm e}^{2}c^{4}/(\sigma_{T}B)}\approx 40$ TeV. Even if the acceleration efficiency is at its extreme maximum $\eta=1$ (Bohm diffusion limit), the attainable energy is only 130 TeV. Local magnetic field strength $B(\phi)$ varies substantially along the orbit—due to changing orbital separation, shock compression, Be-disk crossings, and magnetic reconnection/turbulent amplification—so adopting an orbit-averaged $B\approx 0.2$ G [33] may significantly misestimate the local $E_{\rm max}$. Consequently, despite the detection of photons up to >>200 TeV, a purely leptonic scenario cannot yet be ruled out.
@@
Hadronic processes are expected to be most efficient near periastron, as indicated by the flux enhancement around this orbital phase observed by KM2A. At these phases, the compact object encounters higher ambient densities, for instance through interaction with the circumstellar disk (possibly at its outer edge). Adopting the currently accepted orbital parameters of LS I +61 303 (orbital period $P\approx 26.5$ days, eccentricity $e\approx 0.54$, total mass $M_{\rm tot}\approx 14M_{\odot}$ [24] and then semi-major axis $a\approx 0.4$ au), the radial density profile of the stellar wind can be expressed as $n(r)=\dot{M}/(4\pi r^{2}v_{w}m_{p})$, where we adopt canonical values for the mass-loss rate and velocity, $\dot{M}\sim 5\times 10^{-8}M_{\odot}~{\rm yr}^{-1}$ and $v_{w}=10^{8}~{\rm cm}~{\rm s}^{-1}$. At periastron ($r=r_{p}$), this yields $n(r_{p})\approx 1.9\times 10^{8}~{\rm cm}^{-3}$. Within the Be-star disk, however, densities are 2–4 orders of magnitude higher; we adopt $n_{\rm disk}\sim 10^{11}~{\rm cm}^{-3}$ as a fiducial value. The proton–proton cooling timescale is $t_{pp}=(n\kappa_{pp}\sigma_{pp}c)^{-1}$, with $\sigma_{pp}\approx 40$ mbarn and $\kappa_{pp}\approx 0.5$. In the absence of disk-crossing events, one obtains $t_{pp}\sim 100$ days, which is much longer than the orbital period and thus renders this scenario unlikely, whereas during disk crossing $t_{pp}\sim 4.6$ hours. The interaction efficiency can be estimated as $f_{pp}=t_{\rm esc}/(t_{pp}+t_{\rm esc})$, where the escape/advection time is given by $t_{\rm esc}\sim l/v$, with $l\sim 10^{12}$ cm for a compact interaction zone and $v\sim 10^{9}~{\rm cm}~{\rm s}^{-1}$, leading to $t_{\rm esc}\sim 10^{3}$ s. Consequently, $f_{pp}\approx 0.08$ during disk-crossing episodes and $f_{pp}\approx 1\times 10^{-4}$ otherwise. Since only about one third of the pion energy is transferred into $\pi^{0}\to\gamma\gamma$, the gamma-ray luminosity is related to the proton power by $L_{\gamma}\approx f_{pp}L_{p}/3$. From the KM2A spectrum, we infer a flux of $F_{\gamma,{\rm 25-100TeV}}=1.05\times 10^{-13}~{\rm erg}~{\rm cm}^{-2}~{\rm s}^{-1}$. Adopting a source distance of $d\approx 2.0$ kpc, this corresponds to a gamma-ray luminosity of $L_{\gamma}\approx 5\times 10^{31}~{\rm erg}~{\rm s}^{-1}$. The required proton power is therefore $L_{p}\approx 3L_{\gamma}/f_{pp}$. During disk-crossing phases, this yields $L_{p}\approx 1.9\times 10^{33}$ erg s-1, which is compatible with the spin-down power of a putative pulsar (assuming a pulsar with $P\simeq 270$ ms [7], moment of inertia $I\approx 10^{45}~{\rm g}~{\rm cm}^{2}$, and period derivative $\dot{P}=3\times 10^{-15}~{\rm s}~{\rm s}^{-1}$, the spin-down luminosity is $\dot{E}=4\pi^{2}I\dot{P}/P^{3}\approx 6\times 10^{33}~{\rm erg}~{\rm s}^{-1}$). Conversely, in the non-crossing scenario one obtains $L_{p}\approx 1.5\times 10^{36}~{\rm erg}~{\rm s}^{-1}$, which significantly exceeds the available spin-down power.
@@
Although the observed energy-dependent orbital modulation is only moderately significant, it nevertheless offers valuable insights into the physical conditions across the orbit. As discussed above, the flux enhancement observed by KM2A near periastron may originate from hadronic processes. In the WCDA energy band, however, the flux peak occurs at $\phi\sim 0.6$ (consistent with the VERITAS results), but the light curve exhibits a more complex profile, suggesting the presence of multiple gamma-ray production mechanisms. A hint of UHE-photon concentration around apastron further supports a mixed two-zone scenario. In principle, electrons can be accelerated beyond 100 TeV if the acceleration site is slightly displaced from the binary system. Nevertheless, even if an extended outer accelerator (the Coriolis-induced termination shock [34]) is present at all orbital phases, its ability to produce observable >>100 TeV photons may still be phase-dependent: near periastron the outer shock may be pushed inward and/or compressed by the stronger ram pressure, which can amplify the local magnetic field and shorten synchrotron loss times, thereby reducing the synchrotron-limited electron maximum energy; additionally, the higher local photon and matter densities increase $\gamma\gamma$ opacity and promote electromagnetic cascades that reprocess ultra–high-energy photons to lower energies. By contrast, toward apastron the outer shock is expected to be more extended and less affected by strong radiative losses, favoring acceleration and escape of electrons up to tens or hundreds of TeV and thus enhancing the visibility of >>100 TeV photons. Given the low statistics in the >>100 TeV band, these qualitative effects suffice to reconcile the apparent concentration of the handful of ultra–high-energy events at larger separations, while detailed quantitative modeling must await higher-significance data. Future observations with LHAASO, augmented by coordinated multi-wavelength campaigns, will provide a more comprehensive dataset: phase-resolved spectroscopy at the highest energies, deeper searches for >>100 TeV photons, and contemporaneous X-ray and radio monitoring (to track synchrotron cooling and shock activity) will be particularly valuable. Combined phase-resolved radiative-transfer and cascade modeling, constrained by these multi-band data, should be able to quantify the relative roles of inner/outer accelerators and test the scenario outlined above.
@@
Acknowledgements We would like to thank all staff members who work at the LHAASO site above 4400 meters above sea level year round to maintain the detector and keep the water recycling system, electricity power supply and other components of the experiment operating smoothly. We are grateful to the Chengdu Management Committee of Tianfu New Area for the constant financial support for research with LHAASO data. We appreciate the computing and data service support provided by the National High Energy Physics Data Center for the data analysis in this paper. This research work is supported by the following grants: The National Natural Science Foundation of China(NSFC) No.12522510, No.12393851, No.12393852, No.12393853, No.12393854, No.12205314, No.12105301, No.12305120, No.12261160362, No.12105294, No.U1931201, No.12375107, No.12273038, CAS Project for Young Scientists in Basic Research(No. YSBR-061), Youth Innovation Promotion Association CAS(No.2022010, No.2023275) and in Thailand by the National Science and Technology Development Agency (NSTDA) and the National Research Council of Thailand (NRCT) under the High-Potential Research Team Grant Program (N42A650868).

### v3

@@ I Introduction
Binary systems are prominent sources in the high-energy sky above a few keV, with X-ray binaries being the dominant population [27]. However, only a small number of these systems have been detected at more energetic gamma-ray bands [23, 19]. Among them, a distinct class known as gamma-ray binaries has been identified, characterized by persistent gamma-ray emission and spectral energy distributions ($\nu F_{\nu}$) that peak in the gamma-ray range [23, 20]. All known gamma-ray binaries consist of a compact object in orbit around a massive, young O- or Be-type star, with clear orbital modulation observed across multiple wavelengths [23]. These observed modulations indicate that the gamma-ray emission is likely powered by intra-binary interactions — such as collisions between a relativistic outflow from the compact object and the dense stellar wind or radiation field of the companion star. TeV emission was detected from these compact systems, but it is unclear about the maximal energy reachable by binary systems and whether they are capable of accelerating protons to PeV energies. Probing these systems at ultra-high-energy (UHE; >>100 TeV) range opens a unique window to study extreme particle acceleration, relativistic outflows, and the accretion–ejection processes operating in compact binary environments.
LS I +61 303 is a prototypical gamma-ray binary system composed of a massive Be-type star and a compact object of uncertain nature [41, 26]. Recent report of transient radio pulsations from the direction of LS I +61 303 strongly suggests that the compact object is a rotating neutron star [47]. Located at a distance of $\sim$2 kiloparsecs [43], LS I +61 303 exhibits a well-established orbital period of 26.496(8) days [45], confirmed across optical [39], X-ray [32, 34], and gamma-ray [1, 50] bands. In 2006, the MAGIC telescope detected variable TeV gamma-ray emission modulated by the orbital period, with the short variability time scale suggesting a compact emission origin [9]. This orbital modulation was subsequently confirmed by VERITAS [4]. Both instruments detected enhanced TeV emission near apastron. Besides the well-established orbital modulation, LS I +61 303 is also known to exhibit significant superorbital variability on timescales of $\sim 4.5$ years [8], as well as stochastic TeV flares [12].
The Large High Altitude Air Shower Observatory (LHAASO) is a large hybrid observatory designed to explore the gamma-ray sky from $\sim 100$ GeV to several PeV [5]. It consists of three major detector arrays: the Kilometer Square Array (KM2A), the Water Cherenkov Detector Array (WCDA) and the Wide Field-of-View Cherenkov Telescope Array (WFCTA) [6, 15, 7]. In this paper, we present LHAASO observations of LS I +61 303, spanning roughly 1 TeV to several hundred TeV. These results provide unprecedented access to the UHE regime and offer new constraints on the nature of particle acceleration in gamma-ray binaries.
@@ II Methods
Data from both WCDA and KM2A are used and analyzed independently in this work. The WCDA data are collected with the full array from March 2021 to July 2024, using the number of triggered detectors ($N_{\rm{hit}}$) as a proxy for the energy estimate. The KM2A data used here were collected from December 2019 to July 2024, covering three phases of the array. For KM2A, the primary energy is reconstructed based on the particle density at a fixed distance of 50 m from the shower axis. The data are binned with a resolution of five bins per energy decade. The residual cosmic-ray background after $\gamma$/hadron separation is estimated using the ‘direct integration method’ for both arrays [25]. (Further details regarding the LHAASO detector and event reconstruction are provided in the Supplementary Material and previous performance studies [6, 7].)
The galactic diffuse emission(GDE), produced by the interaction of the cosmic ray ‘sea’ with interstellar gas, may also influence the results [14]. The GDE is assumed to follow the spatial distribution of the Planck dust optical depth map, while its spectrum is modeled by a power-law function. In our joint likelihood fit, both the normalization and the spectral index of the GDE are treated as free parameters. A circular region with a radius of $4^{\circ}$ around LS I +61 303 is chosen as the Region of Interest (ROI). A 3D likelihood fitting process is employed in this analysis, in which the morphological and spectral parameters are fitted simultaneously [16]. Likelihood ratio tests are performed for each potential source in the ROI. A source is added to the model if the improvement in the Test Statistic (TS) exceeds 25 (see Supplementary Material for details).
The sky map in the celestial coordinates is divided into $0.1^{\circ}\times 0.1^{\circ}$ pixels. The TS value at each pixel is calculated using the same likelihood fitting method described above. Specifically, the null hypothesis consists of the background-only model. In the alternative hypothesis, in addition to the background, a point source with its spectral index fixed to the best-fit value is added at the position of each pixel. Since the normalization is the only free parameter, the significance is estimated as $\sqrt{TS}$ according to Wilks’ theorem.
To maintain consistent detector performance, only data taken with the full array from July 2021 to July 2024 are used for the orbital-variability analysis. The orbital phase of each event is calculated as $\phi=((T-T_{0})\bmod P)/P$, where $T_{0}=43366.275$ MJD is the zero-phase epoch [44], $P=26.496$ d is the orbital period [45], and $\bmod$ denotes the modulo operator. We adopt the widely accepted periastron phase $\phi_{\rm peri}=0.23$ [17, 8], although alternative orbital solutions yield a slightly later phase ($\phi_{\rm peri}=0.275$; [11]). The data are divided into ten phase bins per orbit. A likelihood-ratio test is employed, where the null hypothesis assumes constant flux across all bins, whereas the alternative hypothesis allows the flux to vary independently among them. The flux normalization for each phase bin is fitted by assuming the average photon index derived from the full dataset. Under Wilks’ theorem, the resulting TS follows a $\chi^{2}$ distribution with 9 degrees of freedoms.
Figure 1: Test-statistic maps of the region around LS I +61 303 observed by WCDA (A) and KM2A (B), after subtracting the contributions from nearby sources and Galactic diffuse emission. The energy ranges are 1.5–30.5 TeV for WCDA and >25>25 TeV for KM2A, with a sky map pixel size of $0.1^{\circ}\times 0.1^{\circ}$. The gray solid lines represent the significance contours starting from $2\sigma$ with a step of $2\sigma$ (reaching $8\sigma$ for WCDA and $6\sigma$ for KM2A). Green crosses indicate the optical position of LS I +61 303 [46]. The red and black symbols represent the results for WCDA and KM2A, respectively: ‘x’ marks the best-fit positions of LS I +61 303, while dashed circles illustrate the 95% confidence level positional uncertainties, incorporating both statistical fitting errors and systematic pointing uncertainties. White dashed lines indicate the 68% containment regions (convolved with the PSF) of the extended source 1LHAASO J0249+6022. The PSF for each detector is marked by a double-headed arrow.
@@
Figure 3: Phase-resolved emission of LS I +61 303 across different energy bands. (A,B) Gamma-ray flux measured by WCDA (1.5–30.5 TeV) and KM2A (25–100 TeV) as a function of the orbital phase. Downward arrows represent the 95% confidence level upper limits. (C) Estimated energy of individual photon >100>100 TeV detected by KM2A versus their corresponding orbital phases. For clarity, two orbital cycles are shown. The red and blue dash-dotted vertical lines indicate the periastron and apastron phases, respectively, according to the orbital solution from [17].
@@ III Results
Benefiting from the broad energy coverage of LHAASO, we probe the orbital modulation across different energy bands. As shown in Fig. 3, the phase-resolved flux measured by WCDA at 1.5–30.5 TeV extends over most phases and peaks near $\phi=0.6$. A likelihood-ratio test rejects the constant-flux hypothesis with a $p$-value of $0.0053$, corresponding to a significance of $\sim 2.6\sigma$ (one-sided hereafter), which indicates potential marginal variability. In the 25–100 TeV range, the emission is more confined to $\phi=0.3$$0.6$, with a $p$-value of $4.65\times 10^{-5}$ ($\sim 3.9\sigma$). Above 100 TeV, due to limited statistics, we split the data into two broad phase bins: 0.2–0.6 and 0.6–0.2. This yields a $2.4\sigma$ hint for modulation. Interestingly, the UHE photons tend to cluster in the $\phi=0.6$$0.2$ range, differing from the distribution observed at lower energies. However, the current significance remains too low to draw a firm conclusion.
@@ IV Discussion and conclusion
In this study, we report the first definite detection of gamma-ray emission up to the UHE range from the gamma-ray binary LS I +61 303 using LHAASO observations. The extension of the spectral energy distribution to energies nearly an order of magnitude higher than previous measurements (e.g., by MAGIC and VERITAS) not only challenges conventional acceleration models but also provides a unique opportunity to probe the balance between particle acceleration and energy losses in such systems.
@@
In the leptonic scenario, high-energy electrons upscatter stellar photons via inverse Compton (IC) scattering. However, in the UHE regime, the IC cross-section enters the Klein–Nishina limit (For a 100 TeV electron, $\gamma\varepsilon\gg m_{e}c^{2}$ with $\varepsilon=5.5$ eV), significantly suppressing the emission efficiency. Furthermore, the emission spectrum is constrained by the maximum electron energy attainable in the binary system, which is governed by the balance between particle acceleration efficiency, cooling timescales, and the physical size of the acceleration region. Early studies [31, 30] have shown that the maximum energy is sensitive to parameters such as the magnetic field strength, the efficiency of the acceleration mechanism, and the ambient photon density. Assuming an acceleration timescale $\tau_{\rm acc}=E/(\eta eBc)\approx 1.1(E/1{\rm TeV})(\eta/0.1)^{-1}(B/1{\rm G})^{-1}$ s with efficiency factor $\eta\approx 0.1$ and a magnetic field $B\approx 0.2$ G [49], and a synchrotron cooling timescale $t_{\rm syn}=6\pi m_{e}^{2}c^{3}/(\sigma_{T}B^{2}E)\approx 400(E/1{\rm TeV})^{-1}(B/1{\rm G})^{-2}$ s, the maximum energy $E_{\rm max}$ can be obtained by equating $\tau_{\rm acc}$ to the synchrotron cooling timescale $t_{\rm syn}$, then $E_{\rm max}=\sqrt{6\pi\eta em_{\rm e}^{2}c^{4}/(\sigma_{T}B)}\approx 40$ TeV. Even if the acceleration efficiency is at its extreme maximum $\eta=1$ (Bohm diffusion limit), the attainable energy is only 130 TeV. The local magnetic field strength $B(\phi)$ varies substantially along the orbit, driven by factors such as changing orbital separation, shock compression, Be-disk crossings, and magnetic reconnection or turbulent amplification. Therefore, adopting an orbit-averaged $B\approx 0.2$ G [49] may significantly misestimate the local $E_{\rm max}$. Consequently, despite the detection of photons up to >>200 TeV, a purely leptonic scenario cannot yet be ruled out.
@@
Hadronic processes are expected to be most efficient near periastron, as suggested by the KM2A flux enhancement around this orbital phase. At these phases, the compact object encounters higher ambient densities, for instance through interaction with the circumstellar disk (see Supplementary Material for detailed estimations and other binary parameter settings discussed hereafter). Adopting the currently accepted orbital parameters of LS I +61 303, the radial density profile of the stellar wind can be expressed as $n(r)=\dot{M}/(4\pi r^{2}v_{w}m_{p})$, yielding a wind density at periastron of $n(r_{\rm p})\approx 1.9\times 10^{8}~{\rm cm}^{-3}$, while the characteristic density in the circumstellar disk at periastron is $n_{\rm disk,peri}\sim 10^{11}~{\rm cm}^{-3}$. The proton–proton cooling timescale is $t_{pp}=(n\kappa_{pp}\sigma_{pp}c)^{-1}$, with $\sigma_{pp}\approx 40$ mbarn and $\kappa_{pp}\approx 0.5$. In the absence of disk-crossing events, one obtains $t_{pp}\sim 100$ days, which is much longer than the orbital period and thus renders this scenario unlikely, whereas during disk crossing, $t_{pp}\sim 4.6$ hours. The interaction efficiency $f_{pp}$ also differs between disk-crossing ($f_{pp}\approx 0.06$) and non-disk-crossing ($f_{pp}\sim 10^{-4}$) episodes. Since only about one third of the pion energy is channeled into $\pi^{0}\to\gamma\gamma$, the gamma-ray luminosity is related to the proton power by $L_{\gamma}\approx f_{pp}L_{p}/3$. From the KM2A spectrum, the measured 25–100 TeV flux corresponds to a gamma-ray luminosity of $L_{\gamma}\approx 5\times 10^{31}~{\rm erg}~{\rm s}^{-1}$ for a distance of $d\approx 2$ kpc. The inferred proton power during disk-crossing phases is $L_{p}\approx 2.5\times 10^{33}$ erg s-1, which requires rather stringent conditions to channel a substantial fraction of the pulsar spin-down power into protons; nevertheless, the required energy budget remains compatible with the spin-down luminosity of a putative pulsar ($\dot{E}\approx 6\times 10^{33}~{\rm erg}~{\rm s}^{-1}$). In contrast, in the non-crossing scenario, the required proton power increases to $L_{p}\approx 1.5\times 10^{36}~{\rm erg}~{\rm s}^{-1}$, which significantly exceeds the available spin-down power. While a purely hadronic origin challenges the available energy budget when considering the broadband emission, a mixed lepto-hadronic scenario naturally accommodates the constraints (see Supplementary Material for details). This scenario is further supported by archival X-ray and TeV observations (see Fig. S4), which exhibit correlated orbital modulation peaking at phase $\phi\sim 0.6$, and suggest a common leptonic origin for the emission at these phases.
@@
Although the observed energy-dependent orbital modulation is only moderately significant, it nevertheless offers valuable insights into the physical conditions across the orbit. While the leptonic model naturally accounts for the WCDA emission via IC scattering, its extension to the UHE regime is challenged by Klein–Nishina suppression and synchrotron cooling limits described above. So we propose a secondary mixed scenario where the dominant particle population shifts with energy. In the 25–100 TeV band, the post-periastron emission peak sustained through inferior conjunction suggests a target-density-driven hadronic enhancement, while also marking a gamma-ray-transparent phase[22, 21]. This interpretation is reinforced by the simultaneous GeV orbital light curve (Fig. S4), which also exhibits a prominent peak near periastron. The synchronization of GeV and multi-TeV signals strongly implies that the compact object encounters a substantially denser medium at this phase, e.g. Be-disk crossing event, which boosts both the inverse Compton emission and the $pp$ interaction rate. In contrast, the >>100 TeV photons clustering toward apastron indicate a transition to an ‘acceleration-driven’ regime. As described by Zabalza et al. (2013) [48], orbital motion drives a Coriolis-terminated shock extending beyond the binary, leading to a weaker-field zone that suppresses synchrotron cooling. Although both leptons and hadrons can in principle be accelerated to UHE energies under these conditions, the Klein–Nishina suppression of IC scattering severely limits leptonic efficiency, leaving hadrons as the only viable candidate for the >>100 TeV emission. This implies that LS I +61 303 operates as a Galactic PeVatron candidate. Regarding the orbital visibility, it is worth noting that the detected photons emerge in a phase window that avoids the strongest $\gamma\gamma$ opacity expected around superior conjunction [22, 21]. Given the current low statistics in the >>100 TeV band, these qualitative effects suffice to reconcile the apparent concentration of the handful of UHE photons at larger separations, while detailed quantitative modeling must await higher-significance data. Future observations with LHAASO, augmented by coordinated multi-wavelength campaigns, will provide a more comprehensive dataset. Specifically, phase-resolved spectroscopy at the highest energies, deeper searches for >>100 TeV photons, and contemporaneous X-ray and radio monitoring (to track synchrotron cooling and shock activity) will be particularly valuable. Combined phase-resolved radiative-transfer and cascade modeling, constrained by these multi-band data, should be able to quantify the relative roles of inner/outer accelerators and test the scenario outlined above.
@@
Acknowledgements We would like to thank all staff members who work at the LHAASO site above 4400 meters above sea level year-round to maintain the detector and keep the water recycling system, electricity power supply, and other components of the experiment operating smoothly. We are grateful to the Chengdu Management Committee of Tianfu New Area for the constant financial support for research with LHAASO data. We appreciate the computing and data service support provided by the National High Energy Physics Data Center for the data analysis in this paper. This research work is supported by the following grants: National Key Research and Development Program of China under grants 2025YFE0202600, The National Natural Science Foundation of China (NSFC) No.12522510, No.12393851, No.12393852, No.12393853, No.12393854, No.12205314, No.12105301, No.12305120, No.12261160362, No.12105294, No.U1931201, No.12375107, No.12273038, CAS Project for Young Scientists in Basic Research (No. YSBR-061), Youth Innovation Promotion Association CAS (No.2022010, No.2023275) and in Thailand by the National Science and Technology Development Agency (NSTDA) and the National Research Council of Thailand (NRCT) under the High-Potential Research Team Grant Program (N42A650868).